Week of 050801

Open Actions from last week:

  • testing of smoke tests (Jamie)
  • new FTS release / deployment plans.
  • Pilot DB reboot Tuesday DONE
  • New SRM install Tuesday DONE
  • SC3 DB reboot Wed DONE
  • Castor2 DB reboot Wed NOT NEEDED
  • Maarten: Look into GRIDVIEW numbers
  • Korean site for CMS joining the challenge. Need to start the process now.=WONTDO=
  • QF for myproxy/oracle issue in FTS (install new RPM + restart). For thursday

On shift:

  • Jamie/Simone

Monday:

Log: Nothing

New Actions:

Tuesday:

Log: ADC got an alarm from Operators, but with no machine name. Once they worked out the name, they restarted the machine(FTS pilot)

New Actions:

  • Oplapro76 has a hardware problem - needs a FS rebuild (Vlado will clean up and co-ordinate)

Discussion:

  • Oracle port of VOMS has been done. Maria back this week.
  • Maite announced 1.3 gLite. this has the inter-vo stuff and includes configuration.
  • QF for myproxy - still waiting for it to be released - it's in the queue.

Wednesday

Log: Scheduled intervention for oracle patch. Restarted LFC/FTS. Messages from operator, but hadn't mentioned to them that we were bringing down the service. Need to check why not called. Had "unscheduled" patch at 17.10. Restarted daemons and all ok. upgraded SRMs - all looks ok. SRM copy now supported- Peter/Gavin to test.

Actions:

  • Integrate R-GMA procedures into second level support wiki (Shift crew) ADDED
  • Write plan for august tuning phase (James) DONE
  • Check why not called on NO_CONTACT alarms (vlado) DONE procedure was ok.
  • Contact T1 sites to migrate castor SRM to v2. (CASTOR Team)
  • Network tests this evening. DONE
  • QF upgrade for FTS patch DONE
  • GRIDVIEW DB to be upgraded (Monday)
  • Check for "lost" files ( Vlado)
  • Jamie - create intervention log DONE
  • Look at new edg-mkgridmap (Maarten/James) DONE
  • FTS Sensors - Scripts to write (Gavin)

Discussion:

Thursday

Log:

Actions:

  • LSF Scheduler plugin replacement (J-d/Vlado)
  • Laurence - R-GMA can give a tutorial
  • FTS Procedures (James)
  • Intervention log: a temporary page has been created, pending adoption of Miguel Anjo's solution. Please log in reverse chronological order with timestamp, intervention text, intervenor (see log for an example).

Discussion:

  • Document on production integration - Patricia to produce document for this morning.
  • Meeting scheduled for next week with FIO/FS on mid range servers for FTS.

Friday

Log: No Problems.

Actions:

  • TODAY do QF myproxy upgrade DONE

Discussion:

  • Rado reported that instance monitoring will be deployed on our DB Nodes.

Edit | Attach | Watch | Print version | History: r8 < r7 < r6 < r5 < r4 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r8 - 2006-02-07 - PeterJones
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback